Calculate the distance travelled by a car with a velocity of 30/s it increase its speed to 60m/s in 2 hours.assume a acceleration is constant
Answer by Alan3354(69443)

Calculate the distance travelled by a car with a velocity of 30/s it increase its speed to 60m/s in 2 hours. assume a acceleration is constant
=================
Avg speed = (30+60)/2 = 45 m/sec
D = 45*2*3600 meters